CONSIDER A MARKET WHERE SUPPLY AND DEMAND ARE GIVEN BY Qxs=-10+Px AND Qxd=56-2Px. SUPPOSE THE GOVERNMET IMPOSES A PRICE FLOOR OF $25, AND AGREES TO PURCHASE AND DISCARD ANY AND ALL UNITS CONSUMERS DO NOT BUY AT THE FLOOR PRICE OF $25 PER UNIT

Determine the cost to the government of buying firms unsold units
Answer #1

Answer:- as = 10 + Px . O d = 56-2Px supplied gelating, quantity demanded a quantity - - 10 +Px = 56-2Px. 3px 256 + 10 & 3 px